Kraft Corporation has a market value equal to its book value. Currently, Kraft has excess cash of $45,000 and other assets of $502,000. Equity is worth $410,000. The firm has 20,500 shares of stock outstanding and net income of $66,000. What will the new earnings per share be if Kraft uses its excess cash to complete a stock repurchase? $2.87 $3.62 $4.70 $3.26 $4.15
Answer #1

Before Repurchase,

Price Per Share = 410,000/20,500

Price per share = $20

Number of shares repurchased = 45000/20 = 2,250 shares

So,

After Repurchase,

Number of Shares outstanding = 20500 - 2250 = 18,250

EPS after repurchase = 66000/18250 = $3.62

So,

Earning per share after repurchase = $3.62

Please "Like" if you find this useful.